OpenStudy (anonymous):

Please help, -3xyz for x = -3/4 , y = 5/9 and z = -8/15

OpenStudy (anonymous):

Its evaluate

OpenStudy (anonymous):

Replace the variable with the numbers they're equal to \[-3\left(- \frac{ 3 }{ 4 } \right)\left( \frac{ 5 }{ 9 } \right)\left(- \frac{ 8 }{ 15 } \right)\] You can write 3 as 3/1 if it makes it easier \[\left(- \frac{ 3 }{ 1 } \right)\left(- \frac{ 3 }{ 4 } \right)\left( \frac{ 5 }{ 9 } \right)\left(- \frac{ 8 }{ 15 } \right)\] From here, you can cancel common factors and then multiply. Or multiply and then reduce at the end

OpenStudy (anonymous):

I did that, but my teacher was like no.

OpenStudy (anonymous):

- 9 / 4?

OpenStudy (anonymous):